OpenStudy (anonymous):

If x is a positive integer such that (x – 5)^2 = 4 and , what is the value of x?

OpenStudy (shamil98):

(x-5)^2 = 4 square both sides. x-5 = 2 add 5 to both sides. x = 7

OpenStudy (solomonzelman):

Done well. @RichGrl20, do you understand why did he set it up as \[(x-5)^2 = 4\] Are you 100% clear with this?

OpenStudy (solomonzelman):

Oh, I am so stupid, sorry!

OpenStudy (anonymous):

\[(x-5)^2 = 4\] \[(x-5) = \pm 4\] solve for x

OpenStudy (solomonzelman):

\(\huge {\color{green} {Whatever,} }\) \(\huge {\color{green} {I'm} }\) \(\huge {\color{green} {out.} }\)

OpenStudy (anonymous):

iits plus minus 2 sorry